Explore the usefulness of using the Human Development Index (HDI) as a key measurement indicator in your country. Explain how can this indicator assist policymakers in decision-making?

In your response you should:

  • Define the HDI and Explain how it is calculated
  • Using examples, determine the usefulness of the indicator
  • Examine some of the areas where this indicator could be used to determine policy initiatives
Answer #1

Human development index(HDI): As an indicator of development was developed by the United Nations Development Programme (UNDP).It is been used in Human Development Report Published since 1989.

Human development Index (HDI) has three indicators of development to measure the level of welfare of the people of the country:

1. Life Expectancy Indicators (LEI)

2. Educational Attainments Indicators (EAI)

3. Standard of Living Indicators (SLI)

1. Life Expectancy Indicators:

It refers to the life expectancy at birth which implies that number of years a newly born baby is expected to live.

2. Educational Attainment:

It refers to the level of education attained by the people of the country on an average basis.It has

(a) Adult literacy rate: It shows the % of people aged 15 and above who can understand read and write.

(b) Gross enrolment ratio: It shows the % of students enrolled at primary, secondary and university level.

3. Standard of Living of the People:

Real GDP per capita is considered as the indicator of the standard of living of the people.  

HDI Example:

HDI can be constructed by first constructing the individual indices of the above three components and then taking the simple average of the indices.

For example:In India the life expectancy at birth is 69 years. Assuming maximum value and minimum value of life expectancy to be 85 and 20 respectively for a particular year, the LEI can be calculated as follows:

LEI = 69-20/85-20 =49/65 = 0.75

EAI can be calculated by dividing no of years of schooling by the expected years of schooling. For india mean years of schooling is 6.4 and expected years of schooling is 12.4

EAI= 6.4/12.4 = 0.52

SLI can be calculated asThe GNI index uses a minimum income of $100 and a maximum of $75,000. The index uses a logarithmic scale.For India percapita based on purchasing power parity is $ 6,353

[log(6,353)-log(100)] / [log(75,000)-log(100)]= 3.80

Now,Geometric mean of these indices gives HDI =

1.482/3 =0.494

The HDI has huge prospects if it has other variables like child nutriton, women health and women empowerment and employment opportunities at sam e time poverty of the nation keeping in mind the down throned people of the society.

It will be of great use if these were used in health sector to increase the life expectancy ,providing education at a affordable to each and everyone, and also providing employment and skill geneartion to increase the percapita income.